OpenStudy (anonymous):

Find the maximum area of a triangle formed in the first quadrant by the x-axis, y-axis and a tangent line to the graph of f=(x + 5)^2

OpenStudy (freckles):

tangent line to y=(x+5)^2 sounds like we need to find y' there

OpenStudy (freckles):

have you done that part yet?

OpenStudy (anonymous):

I found the derivative to be d/dx=2(x+5) but I am unsure where to go from here.

OpenStudy (freckles):

\[y-y_1=m(x-x_1) \\ m=2x_0+5 \text{ and we have } (x_0,y_0) \] hmmm... so looks just pick a random tangent line to y=(x+5)^2  but this particular tangent line I picked you can not draw a triangle with legs on the x and y axis so ...